# Variational Zero-shot Multispectral Pansharpening

### v1

@@ IV-E Analysis of updating $\mathcal{X}$
Table VIII: Results of PS-TV and PSDip on the reduced resolution WV2, WV3 and QB datasets. The best results are in bold
@@
IV-F Discussions of $R(\mathcal{G})$
The proposed PSDip uses a deep neural network to generate the coefficient $\mathcal{G}$, where the network structure is seen as regularization on $\mathcal{G}$. In this section, we consider replacing the deep image prior with other regularization forms in the pansharpening model (6). As mentioned in Sec. III-B, $\mathcal{G}$ contains image structures. The positions of high-frequency and low-frequency parts of $\mathcal{G}$ are consistent with HRMS. Thus, here, we consider regularizing $\mathcal{G}$ by widely-used Total Variation (TV). By specifying $R(\cdot)$ in the model (6) as the TV norm, we derive the following multispectral pansharpening model:
where the specific form of TV norm is
We call problem (16) PS-TV. In this model, $\mathcal{G}$ itself is a variable. Both $\mathcal{X}$ and $\mathcal{G}$ can also be conveniently optimized by iterative updating. In each step, $\mathcal{X}$ and $\mathcal{G}$ are updated about their respective subproblem. Like PSDip, we also used one step of gradient descent to update $\mathcal{X}$ as well as $\mathcal{G}$ and derive the updating forms like (10). The initial value of $\mathcal{G}$ is set as zero.
In Table VIII, we present the results of PS-TV on reduced resolution WV2, WV3 and QB datasets. First, we can clearly see the effectiveness of PS-TV. This further proves that with appropriate regularization on the coefficient $\mathcal{G}$ by representation $\mathcal{X}=\mathcal{G}\odot\hat{P}$, the proposed model (6) is able to solve multispectral pansharpening problem and acquire satisfying results. Second, in most cases, the PSDip performs better and is more stable than PS-TV. Overall, we can see the benefits of regularizing $\mathcal{G}$ by neural network. Specifically, we think the network parameters can be dynamically adjusted during the optimization process to explore the unknown relationship between HRMS and PAN except for the constraints induced by network structure. Except for TV regularization and deep image prior, we believe more regularizations on $\mathcal{G}$ are hopefully to be investigated.

### v2

@@ IV-E Analysis of updating $\mathcal{X}$
IV-F Comparing $\mathcal{G}$ with HRMS
According to Wald Protocol [83, 84], we generate degraded LRMS $\mathcal{Y}_{l}$ and PAN $P_{l}$ from original LRMS $\mathcal{Y}$ and PAN $P$ by blurring and downsampling. The tools we use are provided by [85].The original LRMS $\mathcal{Y}$ is treated as HRMS. Then $\mathcal{G}_{l}$ is calculated by $\mathcal{G}_{l}=\mathcal{Y}\oslash\hat{P}_{l}$, where $\hat{P}_{l}=\hat{P}_{l}^{\prime}+0.01$ and $\hat{P}_{l}^{\prime}$ is generated by histogram-matching from $P_{l}$ and $\mathcal{Y}_{l}$. We use Fourier transform to separate the low and high frequency part from a image. Since $\mathcal{G}$ and HRMS have different physical meanings, we only visually compare them. Fig. 13 shows an example from full resolution QB dataset. Both the reduced LRMS $\mathcal{Y}_{l}$ and $\mathcal{G}_{l}$ are rescaled to [0,1]. We see in this experiment the high-frequency maps and low-frequency maps between $\mathcal{G}_{l}$ and $\mathcal{Y}$ have many similarities. But it seems that the $\mathcal{G}_{l}$ contains more high-frequency components.
Take $\mathcal{Y}$ as HRMS
High frequency
Low frequency
Coefficient $\mathcal{G}_{l}$
High frequency
Low frequency
Figure 13: Compare the high- and low-frequency parts of $\mathcal{G}$ and HRMS.
IV-G Discussions of $R(\mathcal{G})$
The proposed PSDip uses a deep neural network to generate the coefficient $\mathcal{G}$, where the network structure is seen as regularization on $\mathcal{G}$. Thus, different networks can be seen as different $R(\mathcal{G})$s. In this section, we consider two other network as the backbone of our $f_{\theta}$. They are LAGnet [86] and PNN [44]. Quantatitive results on reduced resolution WV2, WV3 and QB datasets are presented in Table VIII. It first shows that changing the backbone network could also produce satisfying fusion results. This implies the effectiveness of deep image prior for the coefficient. Besides, we see that the performance of the three backbone networks has distinctions because they contains different image priors. For WV2 and QB datasets, LAGnet has overall better results. However, it requires much more running time.
Table VIII: Results of different backbone networks on the reduced resolution WV2, WV3 and QB dataset. The backbone networks are LAGnet, PNN, and PanNet. “T” means the running time. The best results are in bold
[…]
Next, we also consider replacing the deep image prior with other regularization forms in the pansharpening model (6). As mentioned in Sec. III-B, $\mathcal{G}$ contains image structures. The positions of high-frequency and low-frequency parts of $\mathcal{G}$ are consistent with HRMS. Thus, here, we consider regularizing $\mathcal{G}$ by widely-used Total Variation (TV). By specifying $R(\cdot)$ in the model (6) as the TV norm, we derive the following multispectral pansharpening model:
where the specific form of TV norm is
We call problem (17) PS-TV. In this model, $\mathcal{G}$ itself is a variable. Both $\mathcal{X}$ and $\mathcal{G}$ can also be conveniently optimized by alternating minimization. In each step, $\mathcal{X}$ and $\mathcal{G}$ are updated about their respective subproblem. Like PSDip, we also used one step of gradient descent to update $\mathcal{X}$ as well as $\mathcal{G}$ and derive the updating forms like (11). The initial value of $\mathcal{G}$ is set as zero.
In Table IX, we present the results of PS-TV on reduced resolution WV2, WV3 and QB datasets. First, we can clearly see the effectiveness of PS-TV. This further proves that with appropriate regularization on the coefficient $\mathcal{G}$ by representation $\mathcal{X}\approx\mathcal{G}\odot\hat{P}$, the proposed model (6) is able to solve multispectral pansharpening problem and acquire satisfying results. Second, in most cases, the PSDip performs better and is more stable than PS-TV. Overall, we can see the benefits of regularizing $\mathcal{G}$ by neural network. Specifically, we think the network parameters can be dynamically adjusted during the optimization process to explore the unknown relationship between HRMS and PAN except for the constraints induced by network structure. Except for TV regularization and deep image prior, we believe more regularizations on $\mathcal{G}$ are hopefully to be investigated.
Table IX: Results of PS-TV and PSDip on the reduced resolution WV2, WV3 and QB datasets. The best results are in bold.
@@ V Conclusion
Appendix A Proof of Theorem 1
Proof.
As long as we find one such $R(\mathcal{G})$ and $\mathcal{G}_{2}$, we will finish the proof.
Let $R(\mathcal{G}):=\|\mathcal{G}-\mathcal{Z}\|_{F}^{2}$. The value of $\mathcal{Z}$ is underdetermined. Then problem (6) becomes
$\min_{\mathcal{X},\mathcal{G}}\leavevmode\nobreak$ $\left\|\mathcal{Y}-(\mathcal{X}\otimes K)\downarrow_{r}\right\|_{F}^{2}+\lambda_{1}\left\|\mathcal{X}-\mathcal{G}\odot\hat{P}\right\|_{F}^{2}+$
$\lambda_{2}\left\|\mathcal{G}-\mathcal{Z}\right\|_{F}^{2}.$ (19)
Let $\mathrm{vec}(\mathcal{X})$ mean flattening a tensor $\mathcal{X}$ into a vector. And define
$\bm{\mathrm{y}}=\mathrm{vec}(\mathcal{Y}),\bm{\mathrm{x}}=\mathrm{vec}(\mathcal{X}),\bm{\mathrm{g}}=\mathrm{vec}(\mathcal{G}),$
$\bm{\mathrm{p}}=\mathrm{vec}(\hat{P}),\bm{\mathrm{z}}=\mathrm{vec}(\mathcal{Z}).$
Then problem (A) can be equivalently written as
$\min_{\bm{\mathrm{x}},\bm{\mathrm{g}}}\leavevmode\nobreak\ \left\|\bm{\mathrm{y}}-A\bm{\mathrm{x}}\right\|_{F}^{2}+\lambda_{1}\left\|\bm{\mathrm{x}}-\bm{\mathrm{g}}\odot\bm{\mathrm{p}}\right\|_{F}^{2}+\lambda_{2}\left\|\bm{\mathrm{g}}-\bm{\mathrm{z}}\right\|_{F}^{2},$ (20)
where $A$ represents downsampling and bluring matrix. It is easily seen that problem (20) has an unique minimum point $(\bm{\mathrm{x}}^{*},\bm{\mathrm{g}}^{*})$. They satisfiy
$\begin{cases}{}&A^{T}(A\bm{\mathrm{x}}^{*}-\bm{\mathrm{y}})+\lambda_{1}(\bm{\mathrm{x}}^{*}-\bm{\mathrm{g}}^{*}\odot\bm{\mathrm{p}})=0,\\ &\lambda_{1}\bm{\mathrm{p}}\odot(\bm{\mathrm{p}}\odot\bm{\mathrm{g}}^{*}-\bm{\mathrm{x}}^{*})+\lambda_{2}(\bm{\mathrm{g}}^{*}-\bm{\mathrm{z}})=0.\end{cases}$
Thus, we can just let
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{Z}=\mathrm{vec}^{-1}\left(\dfrac{B}{\lambda_{1}\lambda_{2}\bm{\mathrm{p}}}\right),$
$\mathrm{where}\leavevmode\nobreak\ B=$ $(\lambda_{1}\bm{\mathrm{p}}^{2}+\lambda_{2})\odot\left[(A^{T}A+\lambda_{1}I)\mathrm{vec}(\mathcal{X}^{*})-A^{T}y\right]$
$-\lambda_{1}^{2}\bm{\mathrm{p}}^{2}\odot\mathrm{vec}(\mathcal{X}^{*})$
Then problem (A) has the minimum point $(\mathcal{X}^{*},\mathcal{G}_{2})$, where $\mathcal{G}_{2}=\mathrm{vec}^{-1}(((A^{T}A+\lambda_{1}I)\mathrm{vec}(\mathcal{X}^{*})-A^{T}y)/(\lambda_{1}\bm{\mathrm{p}}))$. ∎